# Secure numerical simulations using fully homomorphic encryption

### v2

@@ Abstract
Data privacy is a significant concern when using numerical simulations for sensitive information such as medical, financial, or engineering data. This issue becomes especially relevant in untrusted environments like public cloud infrastructures. Fully homomorphic encryption (FHE) offers a promising solution for achieving data privacy by enabling secure computations directly on encrypted data. In this paper, aimed at computational scientists, we explore the viability of FHE-based, privacy-preserving numerical simulations of partial differential equations. We begin with an overview of the CKKS scheme, a widely used FHE method for computations with real numbers. Next, we introduce our Julia-based packages OpenFHE.jl and SecureArithmetic.jl, which wrap the OpenFHE C++ library and provide a convenient interface for secure arithmetic operations. We then evaluate the accuracy and performance of key FHE operations in OpenFHE as a baseline for more complex numerical algorithms. Following that, we demonstrate the application of FHE to scientific computing by implementing two finite difference schemes for the linear advection equation. Finally, we discuss potential challenges and solutions for extending secure numerical simulations to other models and methods. Our results show that cryptographically secure numerical simulations are possible, but that careful consideration must be given to the computational overhead and the numerical errors introduced by using FHE.
@@ 1 Introduction
Partial differential equations (PDEs) are widely used to model phenomena across nearly every scientific field, from physics and engineering to biology and finance. Since many PDEs cannot be solved analytically, numerical methods are commonly used to approximate their solutions. Beyond strictly scientific settings, PDE-based simulations also play a key role in industry and business applications. In certain cases, the input data and simulation results involve highly sensitive information—–such as personal health records, financial details, or proprietary engineering designs–—that must be carefully protected. For instance, finite element simulations of the aorta used in personalized approaches for treating cardiovascular disease rely on patient-specific data [1, 2]. These datasets must be handled securely to comply with regulations like the General Data Protection Regulation (GDPR) in the European Union or the Health Insurance Portability and Accountability Act (HIPAA) in the United States, posing challenges for scientific research as well as practical implementations [3, 4]. Another example arises in population genetics, where the evolution of gene frequencies has long been modeled by reaction-diffusion equations [5, 6, 7]. Due to nonlinear dynamics and complex boundaries, these models are typically not amenable to analytical solutions and are therefore solved numerically [8, 9, 10]. A practical issue here is working with combined datasets from multiple providers (which do not trust each other), while still protecting the privacy of individual genetic information [11, 12, 13].
@@
To safeguard data, encryption is commonly employed both at rest and during transmission. However, once the data is decrypted for processing, it becomes vulnerable to attacks. The data privacy risks associated with decryption for processing become especially pronounced when data must be processed remotely, as is often the case with public cloud computing platforms. Limited local resources or the need for distributed systems may necessitate offloading computations to external cloud infrastructures, thereby exposing decrypted data to potential attacks. With the rapid expansion of cloud computing [14], ensuring data privacy during computation has emerged as a pressing concern for both practitioners and researchers [15, 16].
@@
A possible solution to this conundrum is the use of what is known as fully homomorphic encryption (FHE). With FHE, it is possible to effect operations directly on the encrypted data. That is, the result of an operation will be same as if the data had been decrypted first, processed, and then encrypted again. Since in case of FHE the data is never available in plaintext, even a fully compromised compute system would not leak any sensitive information. The notion of homomorphic encryption (homomorphisms) was introduced by Rivest, Adleman, and Dertouzos, who in 1978 suggested that such a scheme might be possible [17]. The first viable scheme was proposed much later by Gentry in 2009 [18], when he showed that fully homomorphic encryption is achievable using lattice-based cryptography. Since then, several FHE schemes have been proposed, with the Cheon-Kim-Kim-Song (CKKS) scheme being the most popular option for computations with real numbers [19]. Other commonly used schemes include the Brakerski/Fan–Vercauteren (BFV) [20, 21] and Brakerski–Gentry–Vaikuntanathan (BGV) [22] schemes for computations with integers, and the Ducas–Micciancio (DM/FHEW) [23] and Chillotti–Gama–Georgieva–Izabachene (CGGI/TFHE) [24] schemes for evaluating Boolean and small-integer arithmetic circuits. A good overview of FHE schemes and their applications can be found in [16].
@@
While fully homomorphic encryption schemes represent an active field of research, first applications of FHE schemes can be found in various disciplines, such as privacy-preserving machine learning [25], analysis of encrypted medical or genomic data [26, 27, 28, 13], or for processing data in financial services [29]. One area where FHE has not been applied yet is the field of computational physics. We believe that there are mainly three reasons for this. First, the concept of FHE is still relatively new and not widely known outside of the cryptography community. Second, while there are some excellent open-source libraries available for FHE, they still require a comparatively high level of expertise to use, making it difficult to get started for researchers who are not familiar with the intricacies of FHE. Third, due to its limited scope of available operations, implementing numerical algorithms using FHE is not straightforward and requires a good understanding of FHE.
In this paper, we aim to address these issues by examining fully homomorphic encryption from a scientific computing perspective and introducing it to the computational physics community. We begin with a brief overview of the CKKS scheme for computational scientists, and present two FHE packages we developed in the Julia programming language that allow one to easily experiment with novel FHE applications. We then use these packages to analyze the performance and accuracy of basic FHE operations to get a baseline for more complex algorithms. Finally, we will demonstrate the first prototypes of FHE-secured numerical simulations by implementing and evaluating two well-known numerical methods for solving partial differential equations. We will find that while secure numerical simulations through FHE are indeed possible, the computational overhead is currently significant for many practical scenarios, and the numerical errors introduced by the encryption techniques must be carefully considered when designing applications. All code used in this paper, as well as the data generated during the experiments, is available in our reproducibility repository [30].
@@
The remainder of the paper is structured as follows: In Sec. 2, we give an overview of the CKKS scheme for FHE with real numbers. Moreover, we introduce the open-source C++ library OpenFHE and the Julia packages OpenFHE.jl and SecureArithmetic.jl, which provide a convenient interface for secure arithmetic operations in Julia for experimentation and rapid prototyping. We examine the implementation of the CKKS scheme in OpenFHE in Sec. 3, and analyze the performance and accuracy of basic FHE operations. In Sec. 4, we implement several finite difference schemes to solve the linear advection equation using FHE, and in Sec. 5 we analyze the performance and accuracy of the cryptographically secured numerical solvers. Sec. 6 explores challenges and potential solutions for extending secure numerical simulations to other mathematical models and numerical methods. Finally, we summarize our findings in Sec. 7 and provide an outlook on future work.
@@ 2 Fully homomorphic encryption
The fundamental concept of homomorphic encryption is to perform operations on encrypted data without decrypting it first. This is achieved by using a special encryption scheme that allows for the evaluation of arbitrary functions on encrypted data. The outcome of the computation is still encrypted, and needs to be decrypted to obtain the final result. The key property of such an encryption scheme is that the final result is the same as if the encrypted data had first been decrypted, the unencrypted data processed, and the result re-encrypted. Thus, in a sense, a homomorphic encryption scheme is transparent to the application of arbitrary functions to the data. This is illustrated in Fig. 1, where some data $x$ is first homomorphically encrypted, then processed into $f(x)$, and finally decrypted (bottom path). This yields the same result as if the data had been directly processed in the clear (top path).
@@
With partially homomorphic encryption, only one type of arithmetic operation can be performed on the encrypted data, i.e., either addition or multiplication. Somewhat homomorphic encryption allows both addition and multiplication, but only supports the evaluation of a subset of computation circuits. Leveled homomorphic encryption extends this to arbitrary computation circuits of bounded (pre-determined) depth. Finally, fully homomorphic encryption (FHE) enables the evaluation of arbitrary functions of unbounded complexity on encrypted data, and is thus the most powerful type of homomorphic encryption. As we will see later, however, there still exist practical limitations to the maximum complexity of functions that can be evaluated using FHE.
In the following, we will first discuss the basics of the CKKS scheme for the fully homomorphic encryption of real numbers and then introduce the software packages we use in this work. Please note that the descriptions of the CKKS scheme and FHE in general are kept at a high level, geared towards computational scientists who are not necessarily experts in cryptography. A good overview of the fundamental principles underlying FHE can also be found in [31].
@@ 2.1 Secure approximate arithmetic with the CKKS scheme
The CKKS (Cheon-Kim-Kim-Song) scheme [19] is a fully homomorphic encryption method for approximate arithmetic on real numbers, built upon the BGV scheme for FHE with integers [22]. It obtains its quantum-resistant security from the ring learning with errors (RLWE) problem, which is a specialization of the learning with errors problem [32] for polynomial rings over finite fields [33]. Like other FHE methods, CKKS can be used either as a symmetric or asymmetric encryption scheme. In this manuscript, we will solely focus on its use in the asymmetric encryption context, where there is a public key to encrypt data and a separate private key for decryption.
@@
In terms of arithmetic operations that can be evaluated on the ciphertext, the CKKS scheme supports element-wise addition and multiplication with ciphertexts, plaintexts, or scalar values, as well as rotation operations. The rotation operation performs a circular shift of the data elements in the ciphertext. That is, it moves the data elements by a fixed number of positions while preserving their order, with elements that are shifted out of one end being reintroduced at the opposite end. From these three basic operations, more complex functions can be constructed. Each of these operations further adds a small amount of noise to the data. While this is generally harmless for addition and rotation operations, multiplication adds a non-significant amount of noise, which needs to be kept in check through a process known as rescaling. Due to this, and because of how the CKKS scheme is constructed, the multiplicative depth of computations in CKKS, i.e., the number of consecutive multiplications, is limited to a fixed number of levels. To overcome this limitation, a technique called bootstrapping can be applied after all levels of a ciphertext are used. This bootstrapping procedure resets the available multiplicative depth in the ciphertext, enabling more computations. Since bootstrapping may be used repeatedly, it effectively removes the restrictions on the algorithmic depth of computations in CKKS (for many practical applications where approximation error growth is limited).
@@
In the following sections, we will discuss the CKKS scheme and its main operations in more detail. For clarity, scalar quantities (including polynomials) are written in regular font, while vectors are denoted by bold font. While not strictly necessary to understand the remainder of the paper, these sections provide a useful background for the subsequent sections, especially for readers unfamiliar with FHE methods. Good high-level descriptions of the CKKS scheme can also be found in, e.g., [37, 38, 39].
@@ 2.1.1 Plaintext encoding and decoding
For encoding, the input vector is treated as coefficients of a polynomial. This polynomial is first converted to a new representation by evaluating it at complex roots of unity of the $N_{R}$-th cyclotomic polynomial using a procedure similar to an inverse Fast Fourier Transform (FFT). This enables component-wise vector multiplication using FHE. The result of the inverse FFT is then scaled by a large value, the scaling factor (or scaling modulus), and rounded to the nearest integer values to make it compatible with RLWE. Note that at this point, the data is still not not encrypted. Due to the limited precision of (double) floating-point numbers and rounding error, the encoding step already introduces a small error into the data.
@@ 2.1.2 Levels and ciphertext modulus
Therefore, we usually set FHE parameters using the desired multiplicative depth of the computation. We will denote the maximum multiplicative depth as $l_{\text{max}}$. Based on the multiplicative depth, precision requirements, and security level, two main lattice parameters are configured: ring dimension $N_{R}$ (as described in Section 2.1.1) and ciphertext modulus $Q$.
The ciphertext modulus has the highest value at the beginning, and then gets reduced after each multiplication by the rescaling operation, which is conceptually similar to truncation in floating-point arithmetic, and whose purpose is explained later in Sec. 2.1.6. It is thus convenient to use the notation of $Q_{l}$, where $l$ is the current level. We start with $l=l_{\text{max}}$, i.e., $Q=Q_{l_{\text{max}}}$.
@@
Note that for efficiency, the OpenFHE implementation of CKKS (which we use in this paper, see also Sec. 2.2) uses the Residue Number System (RNS) representation of large numbers, with all values $q_{i}$ being co-prime. For the RNS implementation and a fixed ring dimension $N_{R}$, most of FHE operations scale approximately linearly with the level $l$ [40].
@@ 2.1.3 Encryption and decryption
CKKS is a public key encryption scheme with two keys, a public and a private one. The public key $\bm{\mathrm{pk}}$ is used for encryption and can be safely shared, while decryption uses the private key $\bm{\mathrm{sk}}$. The public key is generated using three polynomials: a uniformly random polynomial $a$, a small secret polynomial $s$, and a small error polynomial $e$. The secret key is defined as
@@
$\bm{\mathrm{pk}}\coloneqq(\mathrm{pk}_{0},\mathrm{pk}_{1})=(-a\cdot s+e,a)\pmod{Q_{l}},$ (2)
where $\cdot$ denotes polynomial multiplication. Note that $\mathrm{pk}_{0}\coloneqq-a\cdot s+e$ corresponds to the RLWE problem, which can be informally stated as follows: if $a$ and $\mathrm{pk}_{0}$ are given, it is computationally hard to find $s$ (and $e$). Furthermore, we use $\pmod{Q_{l}}$ to emphasize where arithmetic operations are performed modulo the current ciphertext modulus $Q_{l}$. The error polynomial $e$ is essential for the security of the scheme, as it prevents an attacker from recovering the secret $s$ from the public key $\bm{\mathrm{pk}}$ by solving the corresponding linear systems of equations.
@@
$c_{0}$ $=\mathrm{pk}_{0}\cdot u+e_{1}+\mu=(-a\cdot s+e)\cdot u+e_{1}+\mu\pmod{Q_{l}},$ (3)
@@
[…]
@@
$=\mu+e\cdot u+e_{1}+e_{2}\cdot s=\mu+v\approx\mu,$
where $v=e\cdot u+e_{1}+e_{2}\cdot s$ is small compared to the plaintext message $\mu$ by construction. Thus, the decryption operation is not exact and adds a small error to the decrypted data.

### v3

@@ Abstract
Data privacy is a significant concern when using numerical simulations for sensitive information such as medical, financial, or engineering data—especially in untrusted environments like public cloud infrastructures. Fully homomorphic encryption (FHE) offers a promising solution for achieving data privacy by enabling secure computations directly on encrypted data. Aimed at computational scientists, this work explores the viability of FHE-based, privacy-preserving numerical simulations of partial differential equations. The presented approach utilizes the Cheon-Kim-Kim-Song (CKKS) scheme, a widely used FHE method for approximate arithmetic on real numbers. Two Julia packages are introduced, OpenFHE.jl and SecureArithmetic.jl, which wrap the OpenFHE C++ library to provide a convenient interface for secure arithmetic operations. With these tools, the accuracy and performance of key FHE operations in OpenFHE are evaluated, and implementations of finite difference schemes for solving the linear advection equation with encrypted data are demonstrated. The results show that cryptographically secure numerical simulations are possible, but that careful consideration must be given to the computational overhead and the numerical errors introduced by using FHE. An analysis of the algorithmic restrictions imposed by FHE highlights potential challenges and solutions for extending the approach to other models and methods. While it remains uncertain how broadly the approach can be generalized to more complex algorithms due to CKKS limitations, these findings lay the groundwork for further research on privacy-preserving scientific computing.
@@ 1 Introduction
Partial differential equations (PDEs) are used to model phenomena across scientific fields ranging from physics and engineering to biology and finance. Since many PDEs cannot be solved analytically, numerical methods are used to approximate their solutions in scientific, industry, and business applications. In certain cases, the input data and simulation results involve highly sensitive information—–such as personal health records, financial details, or proprietary engineering designs–—that must be appropriately safeguarded. For instance, finite element simulations of the aorta used in personalized approaches for treating cardiovascular disease rely on patient-specific data [1, 2]. These datasets must be handled securely to comply with regulations like the General Data Protection Regulation (GDPR) in the European Union or the Health Insurance Portability and Accountability Act (HIPAA) in the United States, posing challenges for scientific research as well as practical implementations [3, 4]. Another example arises in population genetics, where the evolution of gene frequencies has long been modeled by reaction-diffusion equations [5, 6, 7] that are solved numerically [8, 9, 10]. A practical issue here is working with combined datasets from multiple providers (which do not trust each other), while protecting the privacy of individual genetic information [11, 12, 13].
@@
To safeguard data, encryption is commonly employed both at rest and during transmission. However, once the data is decrypted for processing, it becomes vulnerable to attacks. The data privacy risks associated with decryption for processing become pronounced when data must be processed remotely, as is often the case with public cloud computing platforms. Limited local resources or the need for distributed systems may necessitate offloading computations to external cloud infrastructures, thereby exposing decrypted data to potential attacks. With the rapid expansion of cloud computing [14], ensuring data privacy during computation has emerged as a pressing concern for both practitioners and researchers [15, 16].
@@
A possible solution to this conundrum is the use of what is known as fully homomorphic encryption (FHE). With FHE, it is possible to perform computations directly on the encrypted data. That is, the result of an operation will be same as if the data had been decrypted first, processed, and then encrypted again. Since in case of FHE the data is never available in plaintext, even a fully compromised compute system would not leak any sensitive information. The notion of homomorphic encryption (homomorphisms) was introduced by Rivest, Adleman, and Dertouzos, who in 1978 suggested that such a scheme might be possible [17]. The first viable scheme was proposed much later by Gentry in 2009 [18], when he showed that fully homomorphic encryption is achievable using lattice-based cryptography. Since then, several FHE schemes have been proposed, with the Cheon-Kim-Kim-Song (CKKS) scheme being the most popular option for computations with real numbers [19]. Other commonly used schemes include the Brakerski/Fan–Vercauteren (BFV) [20, 21] and Brakerski–Gentry–Vaikuntanathan (BGV) [22] schemes for computations with integers, and the Ducas–Micciancio (DM/FHEW) [23] and Chillotti–Gama–Georgieva–Izabachene (CGGI/TFHE) [24] schemes for evaluating Boolean and small-integer arithmetic circuits. A good overview of FHE schemes and their applications can be found in [16].
@@
Fully homomorphic encryption schemes represent an active field of research, and applications of FHE schemes are used in disciplines such as privacy-preserving machine learning [25], analysis of encrypted medical or genomic data [26, 27, 28, 13], or for processing data in financial services [29]. One area where FHE has yet to be applied is the field of computational physics. There are mainly three reasons for this: First, the concept of FHE is still new and not well known outside of the cryptography community. Second, while there are some excellent open-source libraries available for FHE, they currently require a high level of expertise to use. Third, since FHE only supports a very limited subset of the usual arithmetic operations, implementing secure numerical algorithms is not straightforward and requires a good understanding of FHE. In this paper, we aim to address these issues by examining fully homomorphic encryption from a scientific computing perspective and introducing it to the computational physics community.
Two representative scenarios highlight the practical relevance of FHE-secured numerical simulations. Both involve a data owner (holding sensitive model parameters or input data) and a compute provider (offering computational resources and/or proprietary simulation software). In the first scenario, the data owner must ensure the confidentiality of the input data, for example, a medical researcher running patient-specific simulations on a public cloud platform. In the second scenario, both parties require confidentiality: the input data must remain private, and the compute provider must protect intellectual property in the simulation code. One example is the use of proprietary multiphysics solvers to analyze sensitive design data in aerospace or energy applications. In both scenarios, the workflow proceeds as follows: the data owner locally encrypts the input using a homomorphic encryption scheme such as CKKS and transmits the ciphertexts to the compute provider. The provider runs the numerical simulation directly on the encrypted data and returns the encrypted results. Decryption and post-processing are then performed locally by the data owner.
@@
To explore the feasibility of secure simulations in a scientific computing context, this paper is structured as follows: Sec. 2 begins with a brief overview of the CKKS scheme for computational scientists. Moreover, we introduce the open-source C++ library OpenFHE and the Julia packages OpenFHE.jl and SecureArithmetic.jl, which provide a convenient interface for secure arithmetic operations in Julia for rapid experimentation. We then use these packages in Sec. 3 to analyze the performance and accuracy of basic FHE operations to establish a baseline for more complex algorithms. In Sec. 4, we implement several finite difference schemes to solve the linear advection equation using FHE. They serve as prototypes for FHE-secured numerical simulations, and their performance and accuracy are analyzed in Sec. 5. While secure numerical simulations through FHE are feasible, the computational overhead is currently significant for many practical scenarios, and the encryption-induced numerical errors must be taken into account when designing applications. Sec. 6 explores challenges and potential solutions for extending secure numerical simulations to other mathematical models and numerical methods. Finally, we summarize our findings in Sec. 7 and provide an outlook on future work. All code used in this paper, as well as the data generated during the experiments, is available in our reproducibility repository [30].
@@ 2 Fully homomorphic encryption
The fundamental concept of homomorphic encryption is to perform operations on encrypted data without decrypting it first. This is achieved by using a special encryption scheme that allows for the evaluation of arbitrary functions on encrypted data, such that the result—once decrypted—is the same as if the data had been processed in the clear. Thus, in a sense, a homomorphic encryption scheme is transparent to the application of arbitrary functions to the data. This is illustrated in Fig. 1, where some data $x$ is first homomorphically encrypted, then processed into $f(x)$, and finally decrypted (bottom path). This yields the same result as if the data had been directly processed in the clear (top path).
@@
With partially homomorphic encryption, only one type of arithmetic operation can be performed on the encrypted data, i.e., either addition or multiplication. Leveled homomorphic encryption allows both addition and multiplication, supporting arbitrary computation circuits of bounded (pre-determined) depth. Finally, fully homomorphic encryption (FHE) enables the evaluation of arbitrary functions of unbounded complexity on encrypted data, and is thus the most powerful type of homomorphic encryption.
We begin with a high-level overview of the CKKS scheme for fully homomorphic encryption of real numbers, followed by a brief introduction to the software packages developed for and used in this work. The discussion is intended for computational scientists unfamiliar with cryptography. A more comprehensive treatment of the underlying principles can be found in [31].
@@ 2.1 Secure approximate arithmetic with the CKKS scheme
The CKKS (Cheon-Kim-Kim-Song) scheme [19] is a fully homomorphic encryption method for approximate arithmetic on real numbers, built upon the BGV scheme for FHE with integers [22]. It obtains its security from the ring learning with errors (RLWE) problem, which is a specialization of the learning with errors problem [32] for polynomial rings over finite fields [33]. Like other FHE methods, CKKS can be used either as a symmetric encryption scheme (a single key for encryption and decryption) or an asymmetric encryption scheme (a public key for encryption and a separate private key for decryption). In this manuscript, we solely focus on its use in the asymmetric encryption context.
@@
In terms of arithmetic operations that can be evaluated on the ciphertext, the CKKS scheme supports element-wise addition and multiplication with ciphertexts, plaintexts, or scalar values, as well as rotation operations. The rotation operation performs a circular shift of the data elements in the ciphertext. That is, it moves the data elements by a fixed number of positions while preserving their order, with elements that are shifted out of one end being reintroduced at the opposite end. Rotations are essential for implementing higher-level operations such as vector summation or matrix-vector products. From these three basic operations, more complex functions can be constructed. Each of these operations introduces a small amount of noise into the ciphertext, which slightly degrades the accuracy of the underlying plaintext data. Moreover, multiplication substantially increases the internal noise level of the ciphertext, to the point where it can eventually prevent correct decryption. To avoid this, a process known as rescaling is applied to keep the noise level in check. Due to this, and because of how the CKKS scheme is constructed, the multiplicative depth of computations in CKKS, i.e., the number of consecutive multiplications, is limited to a fixed number of levels. To overcome this limitation, a technique called bootstrapping can be applied after all levels of a ciphertext are used. This bootstrapping procedure resets the available multiplicative depth in the ciphertext, enabling more computations. Since bootstrapping may be used repeatedly, it effectively removes the restrictions on the algorithmic depth of computations in CKKS (for many practical applications where approximation error growth is limited).
@@
In the following sections, we discuss the CKKS scheme and its main operations in more detail. For clarity, scalar quantities (including polynomials) are written in regular font, while vectors are denoted by bold font. While not strictly necessary to understand the remainder of the paper, these sections provide a useful background for the subsequent sections, especially for readers unfamiliar with FHE methods. Good high-level descriptions of the CKKS scheme can also be found in, for example, [37, 38, 39].
@@ 2.1.1 Plaintext encoding and decoding
For encoding, the input vector is treated as coefficients of a polynomial. This polynomial is first converted to a new representation by evaluating it at complex roots of unity of the $N_{R}$-th cyclotomic polynomial using a procedure resembling an inverse Fast Fourier Transform (FFT). This enables component-wise vector multiplication using FHE. The result of the inverse FFT is then scaled by a large value, the scaling factor (or scaling modulus), and rounded to the nearest integer values to make it compatible with RLWE. At this point, the data is still not not encrypted. Due to the limited precision of (double) floating-point numbers and rounding error, the encoding step already introduces a small error into the data.
@@ 2.1.2 Levels and ciphertext modulus
Therefore, we usually set FHE parameters using the desired multiplicative depth of the computation. We denote the maximum multiplicative depth as $l_{\text{max}}$. Based on the multiplicative depth, precision requirements, and security level, two main lattice parameters are configured: ring dimension $N_{R}$ (as described in Section 2.1.1) and ciphertext modulus $Q$.
The ciphertext modulus gets reduced after each multiplication by the rescaling operation. This operation discards the least significant part of the ciphertext—analogous to truncation in floating-point arithmetic—thereby controlling noise growth and maintaining a fixed scale for subsequent operations. Its purpose is explained later in Sec. 2.1.6. It is thus convenient to use the notation of $Q_{l}$, where $l$ is the current level. We start with $l=l_{\text{max}}$, i.e., $Q=Q_{l_{\text{max}}}$.
@@
For efficiency, the OpenFHE implementation of CKKS (which we use in this paper, see also Sec. 2.2) uses the Residue Number System (RNS) representation of large numbers, with all values $q_{i}$ being co-prime. For the RNS implementation and a fixed ring dimension $N_{R}$, most of FHE operations scale approximately linearly with the level $l$ [40].
@@ 2.1.3 Encryption and decryption
CKKS is a public key encryption scheme with two keys, a public and a private one. The public key $\bm{\mathrm{pk}}$ is used for encryption and can be safely shared, while decryption uses the private key $\bm{\mathrm{sk}}$. The public key is generated using three polynomials: a uniformly random polynomial $a$, a small secret polynomial $s$, and a small error polynomial $e_{\mathrm{pk}}$. The secret key is defined as
@@
$\bm{\mathrm{pk}}\coloneqq(\mathrm{pk}_{0},\mathrm{pk}_{1})=(-a\cdot s+e_{\mathrm{pk}},a)\pmod{Q_{l}}$, (2)
where $\cdot$ denotes polynomial multiplication. The term $\mathrm{pk}_{0}\coloneqq-a\cdot s+e_{\mathrm{pk}}$ corresponds to the RLWE problem, which can be informally stated as follows: if $a$ and $\mathrm{pk}_{0}$ are given, it is computationally hard to find $s$ (and $e_{\mathrm{pk}}$). Furthermore, we use $\pmod{Q_{l}}$ to emphasize where arithmetic operations are performed modulo the current ciphertext modulus $Q_{l}$. The error polynomial $e_{\mathrm{pk}}$ is essential for the security of the scheme, as it prevents an attacker from recovering the secret $s$ from the public key $\bm{\mathrm{pk}}$ by solving the corresponding linear systems of equations.
@@
$c_{0}$ $=\mathrm{pk}_{0}\cdot u+e_{1}+\mu=(-a\cdot s+e_{\mathrm{pk}})\cdot u+e_{1}+\mu\pmod{Q_{l}}$, (3)
@@
To decrypt the encrypted data $\bm{c}=(c_{0},c_{1})$, the private key $\bm{\mathrm{sk}}$ is used to obtain an approximate value of the plaintext message $\tilde{\mu}$ by computing the inner product $\langle\cdot,\cdot\rangle$ of the ciphertext $\bm{c}$ with $\bm{\mathrm{sk}}$ as
@@
$=\mu+e_{\mathrm{pk}}\cdot u+e_{1}+e_{2}\cdot s=\mu+v\approx\mu,$
where $v=e_{\mathrm{pk}}\cdot u+e_{1}+e_{2}\cdot s$ is small compared to the plaintext message $\mu$ by construction. Thus, the decryption operation is not exact and adds a small error to the decrypted data.